9x+3y= -39
-9x+y = 71
y= 71+9x
9x+3(71+9x)= -39
9x+213+27x = -39
x= ?
I know that the coefficient of x is 18 but then I forget how to solve for x

Answer by Alan3354(69443)

9x+3y= -39
-9x+y = 71
y= 71+9x
9x+3(71+9x)= -39
9x+213+27x = -39
----------------
36x + 213 = -39
36x = -252
x = -7
-------
y = 71+9x = 71 - 63
y = 8
---------
***************
Elimination would have been easier for this one, since the coefficients of x are the same.
9x+3y= -39
-9x+y = 71
--------------
0x + 4y = 32
y = 8
----- etc